OpenStudy (anonymous):

Marilyn invested $5000 for one year, at 10% annual interest. Bert also invested $5000 for one year. His interest for the year was $125 more than Marilyn's interest for the year. What was his annual rate of interest?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Marilyn's interest is 10% of $5,000 = $500

OpenStudy (skullpatrol):

Bert's interest is $125 + $500 = $625. What percent is 625 of 5,000?

OpenStudy (skullpatrol):

$$\large \frac{x}{100}=\frac{625}{5,000}$$

OpenStudy (anonymous):

12.5% ?\[\frac{ x }{ 100 } =\frac{ 625 }{ 5000 } 5000x=625(100)\]

OpenStudy (skullpatrol):

Yes, 12.5% is the answer :)

OpenStudy (anonymous):

\[x=125\]

OpenStudy (anonymous):

Thanks you're a big help!

OpenStudy (skullpatrol):

Thanks for asking :)